### v1

Abstract
Foundation models trained with self-supervised learning (SSL) on large-scale histological images have significantly accelerated the development of computational pathology. These models can serve as backbones for region-of-interest (ROI) image analysis or patch-level feature extractors in whole-slide images (WSIs) based on multiple instance learning (MIL). Existing pathology foundation models (PFMs) are typically pre-trained on Hematoxylin-Eosin (H&E) stained pathology images. However, images with special stains, such as immunohistochemistry, are also frequently used in clinical practice. PFMs pre-trained mainly on H&E-stained images may be limited in clinical applications involving special stains. To address this issue, we propose StainNet, a specialized foundation model for special stains based on the vision transformer (ViT) architecture. StainNet adopts a self-distillation SSL approach and is trained on over 1.4 million patch images cropping from 20,231 publicly available special staining WSIs in the HISTAI database. To evaluate StainNet, we conduct experiments on an in-house slide-level liver malignancy classification task and two public ROI-level datasets to demonstrate its strong ability. We also perform few-ratio learning and retrieval evaluations, and compare StainNet with recently larger PFMs to further highlight its strengths. We have released the StainNet model weights at: https://huggingface.co/JWonderLand/StainNet.
keywords:
Special Staining, foundation model, computational pathology.1 Introduction
Modern computational pathology involves the use of deep learning models, such as convolutional neural networks [he2016deep] and ViT [dosovitskiy2020image], for clinical tasks on histopathological images, including cancer screening [bejnordi2017diagnostic], tumor analysis [lu2021ai], morphological retrieval [chen2022fast], and survival prognosis [chen2021whole]. In recent years, SSL methods have gained attention and have been proven to be highly effective [kang2023benchmarking]. These methods design pretext tasks, such as contrastive learning [chen2020simple] and image reconstruction [he2022masked], allowing the model to learn effective representations from unlabeled data and obtain good initialization weights, which can then be directly used as foundation models for pathology image feature extraction.
Most existing PFMs are pre-trained on H&E-stained histological images, which are routinely used and easily accessible in clinical workflows. As a result, these PFMs typically perform well with H&E-stained images. For example, by directly analyzing expert-annotated H&E-stained ROIs, PFMs can be used to identify tumor malignancy or microenvironment patterns [lin2025tumor]. Additionally, when H&E-stained resection or biopsy WSI are cropped into patches, PFMs can extract these into high-dimensional features, which are used in weakly-supervised learning tasks based on MIL [xu2025multiple], such as cancer screening or biomarker detection [campanella2025real]. However, there are other staining techniques in clinical practice. For example, by using antibodies that bind to proteins in tissue cells, immunohistochemistry (IHC) slides are widely used to visualize the presence, location, and expression levels of specific proteins, which are commonly utilized for precision cancer diagnosis, disease subtyping, and drug selection. Due to significant differences in nuclear color, tissue patterns, and texture compared to H&E staining, directly transferring existing PFMs to special-stained images for representation may be limited in clinical performance [li2024unlocking].
To address this issue, we propose StainNet, a ViT-based PFM specifically pre-trained on special-staining images. Specifically, to train StainNet, we first select 20,231 non-H&E-stained WSIs from the publicly available large-scale HISTAI database [nechaev2025histai]. For each WSI, we randomly crop up to 100 tissue-containing patch images to create our pre-training dataset. Next, to develop StainNet, we utilize DINO [caron2021emerging], a lightweight self-supervised learning framework specifically designed for ViT models, enabling StainNet to learn task-agnostic morphological information from a large collection of unlabeled special staining images. Finally, to evaluate StainNet, we construct downstream task datasets at both the ROI and WSI levels and compare its performance with existing pathology foundation models. Through this work, we aim to explore the adaptability of PFMs in clinical tasks involving special-stained images and call for the broader exploration of computational pathology in more diverse clinical pathology image datasets.
2 Related works
2.1 Self-supervised PFMs
Foundation models are accelerating the advancement of computational pathology. By performing SSL on large-scale unlabeled pathology images, models can obtain histology-specific weights, which significantly improve their performance on downstream pathology tasks. For instance, CTransPath [wang2022transformer] uses a MoCov3-based SSL approach [chen2021empirical] and is trained on 15.6 million image patches from 25 anatomical sites to demonstrate its ability in H&E staining image retrieval. HIPT [chen2022scaling] and Lunit [kang2023benchmarking] adopt DINO [caron2021emerging], validating the feasibility of pre-training ViT models on histological images. Given the easy scalability of ViT parameters, recent works have aimed to train on larger datasets with larger ViT models. Examples include UNI [chen2024towards], Virchow [vorontsov2024foundation], and Prov-Gigapath [xu2024whole], which aim to diversify the pre-training dataset with more varied pathology images and leverage more computational resources to build stronger general pathology representation models. Additionally, some studies explore SSL in conjunction with special staining, such as PathoDuet [hua2024pathoduet], which utilizes registered H&E and IHC images to help the model learn cross-staining features, aiding the transfer to more advanced nuclear information.
2.2 Special staining in clinical pathology
H&E staining is the most commonly used histological staining method, providing basic structural information by staining the cell nuclei and cytoplasm. However, H&E staining does not fully reveal certain cell features or tissue details, especially in the detection of immune responses or protein expression. As a result, several special staining methods have been developed to address these limitations [gridley1957manual]. Common special staining methods, such as IHC, Masson’s trichrome, and acid phosphatase staining, can highlight specific proteins, cell subtypes, or extracellular matrix components. For example, IHC utilizes antibodies to bind to target proteins, making it a valuable tool for cancer diagnosis and disease subtyping [magaki2018introduction]; Masson’s trichrome staining is employed to visualize elastic fibers in tissues, particularly in fibrosis-related research [lefkowitch2006special]. Recently, the development of computational pathology has accelerated research into the application of computer vision models to these special staining images, such as lymphocyte detection in IHC images [swiderska2019learning] and the use of multiple special stains for kidney biopsy evaluation [jayapandian2021development]. Some studies have also explored the direct use of generative models to convert H&E images into special stains to improve diagnostic accuracy [de2021deep, bai2023deep, pati2024accelerating, yan2023unpaired]. However, these special staining images differ significantly from H&E staining in terms of staining mechanisms, texture features, and visual appearance, making it difficult for large-scale pre-trained models based on general H&E staining images to capture the specific features of special staining images [li2024unlocking].
3 Methodology
We develop and evaluate our StainNet through data collection, model pre-training, and downstream testing. The entire process is illustrated in Figure 1.
3.1 Data collection
We use the publicly available large-scale WSI database HISTAI [nechaev2025histai] as our pre-training data. The original HISTAI dataset contains 112,801 WSIs, of which 20,265 are special stains. We first apply the OTSU thresholding algorithm [otsu1975threshold] to exclude 34 special staining WSIs that can not extract foreground tissue, resulting in 20,231 WSIs. Then, for each WSI, we perform non-overlapping pixel sliding window operations on the foreground tissue, and randomly crop 100 image patches (if a WSI contains fewer than 100 patches, all the patches will be selected). In total, we obtain 1,418,938 patches for subsequent StainNet pre-training.
3.2 Model pre-training
Our proposed StainNet adopts the SSL method DINO [caron2021emerging] as the pre-training strategy to let the ViT model learn robust feature representations from special-staining pathology images without manual annotations. DINO consists of two networks with the same architecture: a student network optimized via gradient descent and a teacher network updated using an exponential moving average (EMA) of the student weights. We follow the original DINO approach for data augmentation and multi-crop strategies. Specifically, we generate a set of global and local views for each pathology image, and the student takes all views as input, while the teacher only receives the global views. By minimizing the cross-entropy loss between the output probability distributions of the two networks, the student is encouraged to infer global tissue context solely from local information. In addition, centering and sharpening operations are applied to the teacher outputs to prevent mode collapse. Figure 2 illustrates some examples of original images and their augmented versions. The detailed hyperparameters are provided in the implementation.
3.3 Downstream adaptation
Our pre-trained StainNet can be applied to both ROI and WSI analysis for special staining images. For ROI tasks, the extracted features from StainNet can be passed to either a single linear layer (linear probe) or a multilayer perceptron (MLP probe) consisting of two linear layers with a activation function for downstream classification. For WSI tasks, we first crop a set of tissue-containing patch images and then perform offline feature extraction using StainNet. The resulting feature sequence (where is the number of image patches and is the feature dimension) can be fed into a series of MIL models to obtain slide-level representations and classification scores.
4 Experiments
4.1 Downstream datasets and implementation
We evaluate StainNet on one private WSI task NTUH-Ki67-Liver, and two publicly available ROI tasks BCI [liu2022bci] and MIST [li2023adaptive]. The NTUH-Ki67-Liver 3-class dataset consists of 46 Ki67 benign liver slides, 43 Ki67 hepatocellular carcinoma slides, and 50 Ki67 normal liver slides. All slides are collected from the Affiliated Hospital of Nantong University and scanned using the SQS-120P scanning system from Shenzhen Shengqiang Technology Co., Ltd. The annotations are performed and verified at the slide level by two pathologists. The BCI dataset is a collection of 4,870 paired H&E-IHC images showing HER2 biomarker expression. We use the IHC images and categorize them into four HER2 expression levels: 0, 1+, 2+, and 3+, with 240, 1,153, 2,142, and 1,355 samples, respectively. The MIST dataset is a large-scale H&E-IHC paired image dataset. We use the IHC images corresponding to four biomarkers: 5,642 HER2, 5,361 Ki67, 5,153 ER, and 5,139 PR to construct a 4-class dataset for evaluating the ability of PFMs to identify biomarkers across different staining conditions.
We use the AdamW optimizer with a learning rate of to pre-train StainNet for 100 epochs with a batch size of 256. For multi-cropping, global crops are sampled using a random resize scale in the range of 0.4-1.0, and 8 local crops are generated with a scale in the range of 0.05-0.4. For EMA design, we increase teacher momentum linearly from 0.9995 to 1.0 throughout training. The weight decay is scheduled from 0.04 to 0.4, and the teacher temperature is set to 0.04 without warm-up and last layer freezing. We use ViT-Small, pre-trained on ImageNet datasets [deng2009imagenet], as the initial backbone and weights for our StainNet. We also enable normalization of the last layer to stabilize the training process. The experiment is conducted on a single NVIDIA RTX PRO 6000 96GB GPU with mixed-precision and takes approximately 3.04 days to complete. We select the final epoch checkpoint as the StainNet weights used in all subsequent experiments.
For downstream fine-tuning, we use AdamW with a learning rate of and a weight decay of . We set 20 epochs with a batch size of 1 for the WSI task and 20 epochs with a batch size of 128 for the ROI task. Early stopping is set to 5 epochs for both functions. For classification tasks, we save the fine-tuned weights and report accuracy, balanced accuracy, AUC, and F1-score from each epoch that achieves the best balanced accuracy. For retrieval tasks, we use cosine similarity for image matching and report recall@K and mAP@K, where K ranges from 1 to 20. All experiments are conducted using stratified 5-fold cross-validation and full-precision computing on a single NVIDIA RTX 3090 24GB GPU. All baseline methods are trained using the same settings.
| Patch encoder | MIL encoder | ||||||
| Bal acc. | ABMIL | SiMLP | TransMIL | WiKG | AMDMIL | S4MIL | Overall |
| ResNet-50 | 0.7980 | ||||||
| CTransPath | 0.7940 | ||||||
| PathoDuet | 0.6957 | ||||||
| HIPT | 0.7943 | ||||||
| Lunit | 0.8470 | ||||||
| StainNet (Ours) | 0.8769 | ||||||
| Patch encoder | MIL encoder | ||||||
| AUC | ABMIL | SiMLP | TransMIL | WiKG | AMDMIL | S4MIL | Overall |
| ResNet-50 | 0.9019 | ||||||
| CTransPath | 0.9127 | ||||||
| PathoDuet | 0.8382 | ||||||
| HIPT | 0.8964 | ||||||
| Lunit | 0.9275 | ||||||
| StainNet (Ours) | 0.9492 | ||||||
| Patch encoder | MIL encoder | ||||||
| F1-score | ABMIL | SiMLP | TransMIL | WiKG | AMDMIL | S4MIL | Overall |
| ResNet-50 | 0.7931 | ||||||
| CTransPath | 0.7671 | ||||||
| PathoDuet | 0.6668 | ||||||
| HIPT | 0.7801 | ||||||
| Lunit | 0.8409 | ||||||
| StainNet (Ours) | 0.8706 | ||||||
4.2 Comparison with PFMs of similar parameter scale
We compare the proposed StainNet with five PFMs of comparable parameter size: CTransPath [wang2022transformer], PathoDuet [hua2024pathoduet], HIPT [chen2022scaling], Lunit [kang2023benchmarking], and ResNet-50 (ImageNet pre-trained) [he2016deep]. For the WSI task, we evaluate six MIL models: (1) ABMIL [ilse2018attention], a gated-attention MIL,(2) SiMLP [li2025can], a method combining unsupervised average pooling with an MLP classifier, (3) TransMIL [shao2021transmil], a transformer-based MIL, (4) WiKG [li2024dynamic], a dynamic graph representation method, (5) AMDMIL [ling2024agent], a method combining masked denosing with agent tokens, and (6) S4MIL [fillioux2023structured], a state space model-based MIL. For the ROI tasks, we freeze each PFM and apply two probe settings: a linear probe and an MLP probe for fine-tuning. Table 1 and Table 2 respectively show the results on the slide-level NTUH-Ki67-Liver task and ROI-level BCI and MIST tasks.
For NTUH-Ki67-Liver, StainNet consistently achieves the best overall performance across all MIL encoders. We also observe that lightweight MIL encoders exacerbate the performance gap between H&E-stained image pre-trained PFMs, whereas more complex MIL encoders mitigate this gap. For example, StainNet outperforms the second-best PFM in terms of balanced accuracy by 5.4% with ABMIL, whereas the margin narrows to 2.5% when using WiKG. For BCI and MIST, StainNet continues to outperform significantly, especially on MIST, where StainNet surpasses the popular CTransPath by 11.7% and 15.4% in balanced accuracy under the linear and MLP probe settings.
| Patch encoder | Linear probe | MLP probe | ||||||
| BCI 4 classes datasets | Acc. | Bal acc. | AUC | F1-score | Acc. | Bal acc. | AUC | F1-score |
| ResNet-50 | ||||||||
| CTransPath | ||||||||
| PathoDuet | ||||||||
| HIPT | ||||||||
| Lunit | ||||||||
| StainNet (Ours) | ||||||||
| Patch encoder | Linear probe | MLP probe | ||||||
| MIST 4 classes datasets | Acc. | Bal acc. | AUC | F1-score | Acc. | Bal acc. | AUC | F1-score |
| ResNet-50 | ||||||||
| CTransPath | ||||||||
| PathoDuet | ||||||||
| HIPT | ||||||||
| Lunit | ||||||||
| StainNet (Ours) | ||||||||
4.3 Comparison with larger PFMs
We further compare StainNet with three recently proposed larger PFMs: PathOrchestra [yan2025pathorchestra], GPFM [ma2025generalizable], and UNI [chen2024towards] as shown in Figure 3. Although StainNet has nearly 14 fewer parameters than them (22M vs 303M), it still achieves the best performance. For example, StainNet surpasses these three large PFMs by 3.72%, 4.70%, and 11.15% in terms of balanced accuracy (80.09% vs. 76.37%, 75.39%, and 68.94%). These results demonstrate that representations learned from large-scale special-stain images provide substantial advantages over H&E-based PFMs when applied to special-stain tasks.
4.4 Comparison of image retrieval
We also examine the retrieval capability of StainNet to assess its semantic representation quality on special staining histology images. We compare StainNet with CTransPath and UNI on the MIST dataset, as shown in Figure 4. We observe that StainNet consistently achieves the best Recall@K and mAP@K across all choices of K, indicating that it has learned a well-structured embedding space for special staining images.
4.5 Effectiveness of different fine-tuning data ratios
We further evaluate the fine-tuning performance of StainNet under varying amounts of downstream special-stain training data to assess its learning capability. Figure 5 presents the results of StainNet, CTransPath, and UNI on the MIST dataset using linear and MLP probes under different proportions of fine-tuning samples. We observe that although StainNet performs slightly worse than UNI under the 10% data setting with a linear probe, it achieves the best performance in all other settings and substantially outperforms the CTransPath of similar parameter size. Notably, applying an MLP probe significantly increases the performance gap between StainNet and UNI in low-fine-tuning data scenarios.
4.6 Performance variation under different pre-training iterations
Finally, we investigate the performance variation of StainNet under different pre-training iterations and the behavior of its loss curve. We believe these observations can provide valuable insights for the development of SSL methods in computational pathology, where pre-training details such as the expected loss curve and convergence behavior are often underreported. Figure 6 illustrates the pre-training loss curve over 100 epochs and the corresponding linear probe performance on the MIST dataset at six iteration checkpoints. We observe that the loss plateaus between epoch 30 and 50, but with significant performance improvement during this period. After epoch 60, the overall loss begins to decrease again, but with slight changes in performance. Ultimately, the model reaches its optimal state at the final iteration. These findings suggest that the shift in the DINO pre-training loss may serve as a valuable indicator of representation quality. While additional exploration is required to generalize these observations to other settings, we encourage future SSL development in computational pathology to consider allocating more training epochs, which may be crucial for achieving optimal model performance.
5 Conclusion and future works
In this work, we propose StainNet, a lightweight pathology foundation model specifically for special staining histopathology image representation. Utilizing the DINO self-supervised framework, StainNet is trained from over 1.4 million unlabeled images. We evaluate StainNet on one in-house WSI task and two public ROI-level tasks, including classification, few-ratio settings, and image retrieval. The results demonstrate that StainNet learns strong and robust representations for special staining pathology images.
For future work, we plan to develop a larger StainNet by expanding the pre-training dataset, scaling the model parameter size, and evaluating it on broader downstream tasks. We will also explore integrating StainNet with H&E-based PFMs for multimodal applications such as survival prognosis and therapy response prediction. We believe that StainNet will further accelerate the research and applications in computational pathology involving diverse histological modalities.
This work is supported by the National Natural Science Foundation of China (NSFC) (82430062), the Shenzhen Engineering Research Centre (XMHT20230115004), and the Tsinghua Shenzhen International Graduate School Cross-disciplinary Research and Innovation Fund Research Plan (JC2024002). We thank Qiehe Sun for providing GPU support.